OpenStudy (anonymous):

Will Fan & Medal :) 6. Simplify the radical expression. Square root 7 + 3 square root 7 A) 2 square root 7 B) 4 square root 14 C) 4 square root 7 D) 2 square root 14

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \sqrt{7}+3\sqrt{7} }\) like this?

OpenStudy (solomonzelman):

Well, what is \(\color{#000000 }{ \displaystyle x+3x }\) ?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle x+3x=4x }\) \(\color{#000000 }{ \displaystyle z+3z=4z }\) \(\color{#000000 }{ \displaystyle \sqrt{7}+3\sqrt{7}=? }\)

OpenStudy (anonymous):

yea it dos but you dont know if its her

OpenStudy (anonymous):

@Annonomys If I stole that picture from her then why doesn't she have any other pictures or videos?

OpenStudy (anonymous):

he justa meani